A plane electromagnetic wave travels in dielectric medium of relative permittivity 9. Relative to free space, the velocity of propagation in the dielectric is


1
Expert's answer
2021-01-10T18:29:48-0500

The velocity of propagation can be found as follows:


"v=\dfrac{1}{\sqrt{\epsilon\cdot \mu}}=\dfrac{1}{\sqrt{9\cdot 1}}=\dfrac{1}{3}."

Therefore, relative to free space, the velocity of propagation in the dielectric is reduced by a factor of 1/3.

Answer:

Relative to free space, the velocity of propagation in the dielectric is reduced by a factor of 1/3.